# NuLite - Lightweight and Fast Model for Nuclei Instance Segmentation and Classification

### v1

@@ 5 Discussion
In this section, we draw back the discussion of our experimental results. According to the training results in PanNuke, we can assert that our model is equivalent to CellViT-SAM-H and, almost in every analyzed case, better than CellViT-256 in terms of each analyzed metric. Still, we can also assert that our model is less complex than CellViT, especially considering the version with SAM-H with backbone. In fact, the complexity and time inference analysis proved that our model is up about 8 times faster, with parameters up to 40 times lower. Moreover, the comparative analysis of NuLite and CellViT across MoNuSeg, CoNSeP, and GlySAC datasets highlights several key findings related to model performance and generalization capabilities in medical image segmentation tasks. The NuLite models, especially NuLite-H, demonstrate strengths in handling larger input sizes, showing superior or competitive performance against the state-of-the-art CellViT models. This makes them particularly valuable for applications requiring extensive spatial analysis, such as large-scale tissue image segmentation. The consistently high scores across multiple metrics and categories underscore the versatility and robustness of NuLite-H, positioning it as a significant advancement in the field. One notable aspect of the study was comparing performance using different patch sizes (256x256 pixels with 64-pixel overlap vs. 1024x1024 pixels). The results demonstrate that using larger patches (1024x1024 pixels) does not negatively impact the performance and may even slightly improve it in some cases. This is consistent with the findings of [11], which suggest that larger patch sizes can maintain or enhance the accuracy of multi-class metrics without compromising the ability of the model to delineate fine details. The slight performance variations between the two patch sizes across different models indicate that larger patches can be effectively utilized in NuLite and CellViT models, potentially simplifying the preprocessing pipeline and reducing computational overhead. The performance metrics across MoNuSeg, CoNSeP, and GlySAC datasets indicate that CellViT and NuLite are robust in handling diverse data types. However, the NuLite models, especially the medium (NuLite-M) and high (NuLite-H) variants, consistently show competitive or superior performance in several metrics compared to CellViT. Notably, in the MoNuSeg dataset, which focuses solely on segmentation, NuLite-T achieved the highest recall (R_d) of 0.910 with 256x256 patches, underscoring its ability to detect relevant instances accurately. Similarly, NuLite-H demonstrated superior recall and F1 scores in the CoNSeP dataset, which involves more complex tissue classification tasks. The CoNSeP dataset, aligned with PanNuke nuclei types, provided a challenging environment for testing multi-class segmentation capabilities. Here, NuLite-H excelled, particularly in the Neoplastic and Miscellaneous categories, suggesting that the model is adept at handling various tissue types and complex boundaries. The strong performance in the Miscellaneous class, which includes a diverse range of tissues, further underscores the model’s versatility. On the other hand, CellViT-SAM-H showed strong performance in the Epithelial class, indicating its efficacy in distinguishing epithelial tissues with high segmentation quality. The findings from this study have important implications for using these models in whole-slide imaging (WSI) applications. The ability to effectively use larger patches (1024x1024 pixels) could significantly streamline the process of analyzing large WSI data, reducing the need for extensive patch overlap and potentially accelerating the segmentation process. This is particularly relevant in clinical settings with critical time and computational resources.
@@ 6 Conclusion
In this work, we introduced NuLite, a fast and lightweight convolutional neural network for nuclei instance segmentation and classification in H&E stained histopathological images. With its U-Net architecture featuring three decoders for predicting nuclei, horizontal and vertical maps, and nuclei types, drawing inspiration from HoVer-Net, NuLite demonstrates considerable promise. Furthermore, we provided an extensive experimental setting on data not used for training, such as CoNSeP, MoNuSeg, and GlySAC, proving the ability to generalize our model. Therefore, our model demonstrated a state-of-the-art lightweight model in nuclei instance segmentation classification. In some scenarios, it outperforms also CellViT-SAM-H, which is the current SOTA but more complex and heavy than our NuLite. The study reveals that NuLite, especially its medium and high variants, performs on par with or even outperforms current state-of-the-art models like CellViT. Overall, NuLite represents a significant advancement in automated medical diagnostics, offering speed and accuracy that could enhance analysis efficiency in medical contexts. In future work, we will focus on the ability of our model to embed nuclei, as also shown in [11], and we will try to use its ability in cell-graph classification.
@@ Acknowledgments
This study was partially supported by the PNRR MUR project PE0000013-FAIR. We acknowledge the CINECA award (project FVT-NSC) under the ISCRA initiative for the availability of high-performance computing resources and support.

### v2

@@ 5 Discussion
In this section, we draw back the discussion of our experimental results. According to the training results in PanNuke, we can assert that our model is equivalent to CellViT-SAM-H and, almost in every analyzed case, better than CellViT-256 in terms of each analyzed metric. Still, we can also assert that our model is less complex than CellViT, especially considering the version with SAM-H with backbone. The complexity and inference time analysis proved that our model is up to 13 times faster, with parameters up to 58 times lower, with GFLOPS up to 11 times lower, saving up to 12.25 times of memory amount during the inference. Moreover, the comparative analysis of NuLite and CellViT across MoNuSeg, CoNSeP, and GlySAC datasets highlights several key findings related to model performance and generalization capabilities in medical image segmentation tasks. The NuLite models, especially NuLite-H, demonstrate strengths in handling larger input sizes, showing superior or competitive performance against the state-of-the-art CellViT models. These motivations make them particularly valuable for applications requiring extensive spatial analysis, such as large-scale tissue image segmentation. The consistently high scores across multiple metrics and categories underscore the versatility and robustness of NuLite-H, positioning it as a significant advancement in the field. One notable aspect of the study was comparing performance using different patch sizes ($256\times 256$ pixels with 64-pixel overlap vs. $1024\times 1024$ pixels). The results demonstrate that using larger patches ($1024\times 1024$ pixels) does not negatively impact the performance and may even slightly improve it in some cases. This aspect is consistent with the findings of [12], which suggest that larger patch sizes can maintain or enhance the accuracy of multi-class metrics without compromising the ability of the model to delineate fine details. The slight performance variations between the two patch sizes across different models indicate that larger patches can be effectively utilized in NuLite and CellViT models, potentially simplifying the preprocessing pipeline and reducing computational overhead. The performance metrics across MoNuSeg, CoNSeP, and GlySAC datasets indicate that CellViT and NuLite are robust in handling diverse data types. However, the NuLite models, especially the medium (NuLite-M) and high (NuLite-H) variants, consistently show competitive or superior performance in several metrics compared to CellViT. Notably, in the MoNuSeg dataset, which focuses solely on segmentation, NuLite-T achieved the highest recall (R_d) of 0.910 with $256\times 256$ patches, underscoring its ability to detect relevant instances accurately. Similarly, NuLite-H demonstrated superior recall and F1 scores in the CoNSeP dataset, which involves more complex tissue classification tasks. The CoNSeP dataset, aligned with PanNuke nuclei types, provided a challenging environment for testing multi-class segmentation capabilities. Here, NuLite-H excelled, particularly in the Neoplastic and Miscellaneous categories, suggesting that the model is adept at handling various tissue types and complex boundaries. The strong performance in the Miscellaneous class, which includes a diverse range of tissues, further underscores the model’s versatility. On the other hand, CellViT-SAM-H showed strong performance in the Epithelial class, indicating its efficacy in distinguishing epithelial tissues with high segmentation quality. The findings from this study have important implications for using these models in whole-slide imaging (WSI) applications. The ability to effectively use larger patches ($1024\times 1024$ pixels) could significantly streamline the process of analyzing large WSI data, reducing the need for extensive patch overlap and accelerating the segmentation process.
@@ 6 Conclusion
In this work, we introduced NuLite, a fast and lightweight convolutional neural network for nuclei instance segmentation and classification in H&E stained histopathological images. With its U-Net architecture featuring one decoder and three segmentation heads for predicting nuclei, horizontal and vertical maps, and nuclei types, drawing inspiration from HoVer-Net, NuLite demonstrates considerable promise. Furthermore, we provided an extensive experimental setting on data not used for training, such as CoNSeP, MoNuSeg, and GlySAC, proving the ability to generalize our model. Therefore, our model demonstrated a state-of-the-art lightweight model in nuclei instance segmentation classification. In some scenarios, it also outperforms CellViT-SAM-H, the current SOTA, but is more complex and heavy than our NuLite. The study reveals that NuLite, especially its medium and high variants, performs on par with or even outperforms current state-of-the-art models like CellViT. Overall, NuLite represents a significant advancement in automated medical diagnostics, offering speed and accuracy that could enhance analysis efficiency in medical contexts. In future work, we will delve deeper into the capabilities of our model, particularly its ability to embed nuclei, as also shown in [12]. We aim to leverage this ability in cell-graph classification, opening up new possibilities for our model’s application. Furthermore, we are committed to enhancing the entire pipeline for WSI inference.
@@ Acknowledgments
This study was partially supported by the PNRR MUR project PE0000013-FAIR. We also acknowledge the CINECA award (project FVT-NSC) under the ISCRA initiative for the availability of high-performance computing resources and support.